# AuthGlass: Enhancing Voice Authentication on Smart Glasses via Air-Bone Acoustic Features

### v2

@@ 1. Introduction
Nowadays, smart glasses have been rapidly advancing, with commercial products such as Microsoft HoloLens(Mic, 2025) and Ray-Ban Meta Glasses(Met, 2025) demonstrating their potential across diverse domains, including entertainment, healthcare, education, and industrial applications. As smart glasses are closely integrated with the user’s head that allows unique interaction potential, researchers explored numerous application scenarios accessibility support for visually impaired users(Rajendran et al., 2020a; Khan and Prakash, 2017) and context-aware services(Wang et al., 2024; Meyer et al., 2022).
Among various interaction modalities, voice commands have emerged as one of the most widely adopted approaches on smart glasses(Rok, 2025; Met, 2025). Voice interaction is intuitive, hands-free, and well-suited to scenarios where users’ hands are occupied, such as driving(He et al., 2015; Chang et al., 2018) and cooking(Majil et al., 2022; Bouchard et al., 2020). However, as smart glasses are exposed to ambient voices, such interactions are susceptible to potential attacks, leading to severe security risks(Peng et al., 2016). To migrate this issue, various authentication methods has be proposed, including password entry(Islam et al., 2018), iris recognition (Khade et al., 2021; Li and Huang, 2017) and gaze features(Friström et al., 2019). Nevertheless, voice authentication remains the most natural complement to voice command interfaces(Peng et al., 2016), as it does not require additional interaction modalities and preserves the hands-free advantage(Rajendran et al., 2020b).
Although voice authentication has been extensively studied (Kersta, 1962; Wang et al., 2019b; Zhang et al., 2016; Yan et al., 2019), achieving that on smart glasses remains challenging, as only limited work has examined acoustic features tailored to this unique form factor. In contrast, for other wearable platforms such as smartphones and earables (Zhang et al., 2016; Li et al., 2020; Feng et al., 2017; Blue et al., 2018; Meng et al., 2022; Gao et al., 2021), researchers have investigated a wide range of methods to enable voice authentication(Zhang et al., 2016; Huang et al., 2025; Li et al., 2020, 2024b; Meng et al., 2022; Blue et al., 2018). While these approaches have demonstrated competitive accuracy, their direct transfer to smart glasses is non-trivial due to the unique form factor and acoustic constraints of eyewear.
@@ 2.1. Voice Interface for Smart Glasses
As smart glasses were limited with size and shape, voice user interfaces (VUIs) have long been explored for interactions on smart glasses due to its’ naturalness and convenience. Apart from vision-based and gaze based interaction, which rely on heavy and energy consuming sensors, voice interface only requires lightweight microphones for interaction(Palermo et al., 2025), making it widely incorporated by commercial products like Ray-ban Meta(Met, 2025), Xreal One Pro(Xre, 2025), Rokid Glasses(Rok, 2025), etc. More importantly, voice input is suitable for real-time communication and conversation-based AI functions. Ray-ban Meta utilized voice interface to control AI-powered services, while researchers also explored AI potentials on smart glasses via voice interface, especially for visual impaired people(T P et al., 2025; Gamage, 2024; Nithyaa et al., 2023; Rajendran et al., 2020a). Moreover, voice interface offers a ’touchless’ interaction for users and do not interfere with users’ gaze attention, it is especially suitable for emergency medical services(Zhang et al., 2025) and industries (Kim et al., 2016).
As smart glasses increasingly integrate with users’ personal information, securing their voice-based interfaces has become a critical challenge. However, with the rapid advancement of generative AI, voice interface was hindered by various attacking methods (Kamble et al., 2020; Yi et al., 2023; Ergünay et al., 2015). As smart glasses are often exposed to surrounding speech and ambient noise, they are vulnerable to forged commands injected by adversaries(Li et al., 2024a; Chan et al., 2015), leading to potential severe privacy breaches. Prior works(Yan et al., 2022; Li et al., 2023) show that authentication schemes relying solely on acoustic features remain vulnerable to diverse threats, including replay(Yi et al., 2023; Veesa and Singh, 2025) and synthetic speech attacks(Wenger et al., 2021; Zuo et al., 2024), thereby lacking robustness. Most previous works failed to explore usable acoustic features that could be applied to smart glasses. Thus, with this in mind, we aim to comprehensively explore the around-head acoustic features by glasses, hence promoting the design of voice authentication.
@@ 2.2. Authentication Methods on Smart Glasses
To address the security challenges of smart glasses, researchers have explored a wide range of authentication methods across different modalities. Broadly, existing approaches can be categorized into knowledge-based and biometric-based methods (Chang, 2018). Knowledge-based authentication relies on what users know, with passwords being the most common example (Chan et al., 2015; Islam et al., 2018; Findling et al., 2019; Friström et al., 2019; Winkler et al., 2015). While such methods can provide a reasonable level of security, they often impose additional cognitive and interaction burdens on users (Chang, 2018). In contrast, biometric authentication leverages users’ inherent traits (Boshoff and Hancke, 2025), which are difficult to replicate and require no conscious memorization. A variety of biometric-based techniques have been investigated, including iris recognition (Li and Huang, 2017; Khade et al., 2021), gaze movement patterns (Inoue and Murao, 2024; Zhang et al., 2018), facial shape features (Li et al., 2024a; Isobe and Murao, 2021; Lim et al., 2023), and skull characteristics (Shin et al., 2024; Schneegass et al., 2016). Among these, voice has attracted growing interest as a biometric modality (Peng et al., 2016; Huang et al., 2025), since it can be seamlessly integrated into the voice user interfaces (VUIs) that are already central to the interaction paradigm of modern smart glasses.
However, achieving reliable voice authentication on smart glasses remains challenging, as few studies have comprehensively explored acoustic features specifically suited to this form factor. For other wearable platforms such as smartphones and earables (Zhang et al., 2016; Li et al., 2020; Feng et al., 2017; Blue et al., 2018; Meng et al., 2022; Gao et al., 2021), researchers have investigated a wide range of acoustic representations, including time-domain features (Zhang et al., 2016), frequency-domain features (Huang et al., 2025), time–frequency features (Li et al., 2020, 2024b), and sound-field features (Meng et al., 2022; Blue et al., 2018). While these approaches have demonstrated high accuracy, their direct transfer to smart glasses is non-trivial due to the unique form factor and acoustic constraints of eyewear. In particular, smart glasses face challenges such as limited flexibility in microphone placement, variable distance from the user’s mouth, and increased susceptibility to ambient noise. All of these challenges hinder the straightforward adaptation of existing acoustic authentication techniques. Motivated by these challenges, we investigate features derived from both air- and bone-conducted acoustic domains, and evaluate their effectiveness in supporting robust voice authentication for smart glasses.
@@ 2.3. Liveness Detection for Voice Authentication
Beyond distinguishing users from impostors, voice authentication must verify that commands are genuinely produced by the user, a process referred to as liveness detection(Li et al., 2021, 2024b; Akhtar et al., 2014; Shiota et al., 2015; Zhang et al., 2017, 2022b). To achieve this, researchers have proposed various methods, which can be categorized into speaker detection and speech detection. Speaker detection methods aim to determine whether a given sound originates from an artificial source, typically by exploiting features associated with machine-based speech production mechanisms, such as acoustic features(Ahmed et al., 2020; Li et al., 2021) and magnetic features(Chen et al., 2017). However, these methods are tailored to the specific design characteristics of certain artificial speakers, which inherently limits their generalizability. As speech synthesis and voice conversion technologies continue to advance, such approaches may face increasing risks of being circumvented, thereby undermining their robustness in practical deployment.
Speech detection methods, in contrast, exploit features directly associated with the speaking process to ensure that the captured signal is temporally aligned with the user’s actual speech production (Zhang et al., 2022a; Sahidullah et al., 2017; Li et al., 2024b; Zhang et al., 2024). Building on this idea, researchers have explored a range of approaches, including throat vibrations (Sahidullah et al., 2017), bone-conducted vibrations (Li et al., 2024b), facial muscle movements (Zhang et al., 2024), airflow during articulation (Wang et al., 2019a), and pop sounds (Wang et al., 2019b), among others. While these studies have achieved promising results, most of them depend on sensors that either require close contact with the skin (Huang et al., 2025) or must be placed at locations impractical for everyday wearable devices (e.g., the throat) (Sahidullah et al., 2017). Such constraints not only affect user comfort and social acceptability but also hinder seamless integration into mainstream wearables, particularly smart glasses. To address these challenges and enable practical voice authentication on smart glasses, we further investigate the feasibility of leveraging combined acoustic features for liveness detection.
@@ 3. Understanding Air-Bone Acoustic Features by Smart Glasses
Figure 1. (a) Structure of vocal tract, including vocal cords, tongue and other articulatory organs (b) Previous research revealed that phoneme sounds could be localized by method of Time Difference by Arrival (Zhang et al., 2016). (c) The possible acoustic features related to speech production, including vibration conducted by skin and bones, time delay and energy decrease of sound field in spatial dimension.
@@ 3.1. Principle and Possibility
To understand how smart glasses can leverage voice-related features for authentication, it is essential to briefly review the human speech production system, as illustrated in Fig. 1a. Airflow from the lungs excites the vocal folds, the vocal tract then shapes this excitation into an acoustic realization with characteristic amplitudes and spectra (Badin and Serrurier, 2006). After that, through precise control of the tongue and muscles, humans can articulate different phonemes and ultimately form words and sentences. Crucially, speaker-specific anatomy yields distinct propagation paths, resulting in both post-utterance signals and corresponding transmission paths carrying user-unique signatures for authentication.
However, existing approaches are based on features extracted from either the post-utterance voice or the transmission path, limiting comprehensive usage of the voice generation process. Most researchers focus on post-utterance features including airflow (Wang et al., 2019a), pop-sound (Wang et al., 2019b), and knowledge/content-based cues built from the observed speech signal. In contrast, motion-based methods that leverage IMU sensors in smart glasses exploit propagation-induced vibrations of the head/face but have limited access to the linguistic/ phonic content itself—and in smart glasses(Zhang et al., 2024). We therefore ask: how to integrate voice features with transmission-path cues on smart glasses to enable more robust authentication?
In this paper, we found smart glasses offer two potential sensing avenues that jointly encode voice features and how it propagates (Fig. 1b–c). First, air-conduction (air) microphone arrays in the frame enable capturing both the voice and the path-dependent spatial features of the surrounding sound field shaped by head geometry and motion of the articulator (Zhang et al., 2016). A bone/skin-conduction microphone (bone) at a stable contact point (e.g. nose pads) senses the voiceprint plus the propagation-induced vibrations transmitted through tissue and bone (Liu et al., 2018). Thus, both channels inherently provide comprehensive features along the voice generation process, potentially enabling more robust authentication for smartglasses.
@@ 3.2.1. Design and Implementation
To capture the sound field as accurately and comprehensively as possible, we arranged a total of 14 omnidirectional air-conduction (AC) microphones around the user’s head (infineon IM73D122V01(Inf, 2025)) and oriented most of them downward. To maximize generality, our placement strategy spans multiple commercial form factors — including Ray-Ban Meta (nose bridge and legs), Bose Frames (behind the ears). We also integrated microphones on the rims as some prior work also investigated acoustic-based interaction and health sensing at the same locations (Guo et al., 2025b, a).
In order to capture vibration features, we employed bone-conduction microphones (VPU14AA01 BC)(Son, 2025) for data collection, as they provide reliable sensing when in direct contact with the skin. Since smart glasses only make stable contact with the head around the nose and ears, and pilot studies showed that microphones near the ears were easily affected by hair friction, we mounted the bone-conduction microphones on each side of the nose pads to ensure better recording quality. The overall microphone layout is illustrated in Fig. 2a.
To acquire high-fidelity acoustic signals and extract meaningful insights, both bone-conduction and air-conduction microphones were configured in synchronization with a 96 kHz sampling rate. To achieve this, a collection board was used to connect all microphones simultaneously via flexible flat cables (FFCs), as shown in Fig.2a. The collection board was driven by an Orange Pi(Ora, 2025), which also transmitted data to a PC over a network cable using the TCP protocol. Note that a wire bundle was designed to prevent crosstalk between adjacent FFCs and to keep the cables organized, while maintaining user comfort.

### v3

@@ 1. Introduction
Nowadays, smart glasses have been rapidly advancing, with commercial products such as Microsoft HoloLens(MicrosoftHololens) and Ray-Ban Meta Glasses(MetaRayban2025) demonstrating their potential across diverse domains, including entertainment, healthcare, education, and industrial applications. As smart glasses are closely integrated with the user’s head that allows unique interaction potential, researchers explored numerous application scenarios accessibility support for visually impaired users(VisualImpair; khan2017design) and context-aware services(wang2024g; meyer2022u).
Among various interaction modalities, voice commands have emerged as one of the most widely adopted approaches on smart glasses(RokidGlasses2025; MetaRayban2025). Voice interaction is intuitive, hands-free, and well-suited to scenarios where users’ hands are occupied, such as driving(he2015driving; chang2018design) and cooking(majil2022augmented; bouchard2020smart). However, as smart glasses are exposed to ambient voices, such interactions are susceptible to potential attacks, leading to severe security risks(peng2016continuous). To migrate this issue, various authentication methods has be proposed, including password entry(islam2018glasspass), iris recognition (khade2021iris; li2017accurate) and gaze features(fristrom2019free). Nevertheless, voice authentication remains the most natural complement to voice command interfaces(peng2016continuous), as it does not require additional interaction modalities and preserves the hands-free advantage(rajendran2020design).
Although voice authentication has been extensively studied (kersta1962voiceprint; voicepop; zhang2016voicelive; yan2019catcher), achieving that on smart glasses remains challenging, as only limited work has examined acoustic features tailored to this unique form factor. In contrast, for other wearable platforms such as smartphones and earables (zhang2016voicelive; li2020vocalprint; feng2017continuous; blue20182ma; meng2022your; gao2021voice), researchers have investigated a wide range of methods to enable voice authentication(zhang2016voicelive; huang2025eve; li2020vocalprint; li2024boneauth; meng2022your; blue20182ma). While these approaches have demonstrated competitive accuracy, their direct transfer to smart glasses is non-trivial due to the unique form factor and acoustic constraints of eyewear.
@@ 2.1. Voice Interface for Smart Glasses
As smart glasses were limited with size and shape, voice user interfaces (VUIs) have long been explored for interactions on smart glasses due to its’ naturalness and convenience. Apart from vision-based and gaze based interaction, which rely on heavy and energy consuming sensors, voice interface only requires lightweight microphones for interaction(palermo2025advancements), making it widely incorporated by commercial products like Ray-ban Meta(MetaRayban2025), Xreal One Pro(Xreal2025), Rokid Glasses(RokidGlasses2025), etc. More importantly, voice input is suitable for real-time communication and conversation-based AI functions. Ray-ban Meta utilized voice interface to control AI-powered services, while researchers also explored AI potentials on smart glasses via voice interface, especially for visual impaired people(AI-Powered4Speech2text; gamage2024ai; nithyaa2023eye; VisualImpair). Moreover, voice interface offers a ’touchless’ interaction for users and do not interfere with users’ gaze attention, it is especially suitable for emergency medical services(zhang2025touchless) and industries (kim2016augmented).
As smart glasses increasingly integrate with users’ personal information, securing their voice-based interfaces has become a critical challenge. However, with the rapid advancement of generative AI, voice interface was hindered by various attacking methods (kamble2020advances; yi2023audio; ergunay2015vulnerability). As smart glasses are often exposed to surrounding speech and ambient noise, they are vulnerable to forged commands injected by adversaries(li2024sonicid; chan2015glass), leading to potential severe privacy breaches. Prior works(yan2022survey; li2023security) show that authentication schemes relying solely on acoustic features remain vulnerable to diverse threats, including replay(yi2023audio; veesa2025deep) and synthetic speech attacks(wenger2021hello; zuo2024advtts), thereby lacking robustness. Most previous works failed to explore usable acoustic features that could be applied to smart glasses. Thus, with this in mind, we aim to comprehensively explore the around-head acoustic features by glasses, hence promoting the design of voice authentication.
@@ 2.2. Authentication Methods on Smart Glasses
To address the security challenges of smart glasses, researchers have explored a wide range of authentication methods across different modalities. Broadly, existing approaches can be categorized into knowledge-based and biometric-based methods (chang2018two). Knowledge-based authentication relies on what users know, with passwords being the most common example (chan2015glass; islam2018glasspass; findling2019hide; fristrom2019free; winkler2015glass). While such methods can provide a reasonable level of security, they often impose additional cognitive and interaction burdens on users (chang2018two). In contrast, biometric authentication leverages users’ inherent traits (boshoff2025classifications), which are difficult to replicate and require no conscious memorization. A variety of biometric-based techniques have been investigated, including iris recognition (li2017accurate; khade2021iris), gaze movement patterns (inoue2024user; zhang2018continuous), facial shape features (li2024sonicid; isobe2021person; lim2023c), and skull characteristics (shin2024skullid; schneegass2016skullconduct). Among these, voice has attracted growing interest as a biometric modality (peng2016continuous; huang2025eve), since it can be seamlessly integrated into the voice user interfaces (VUIs) that are already central to the interaction paradigm of modern smart glasses.
However, achieving reliable voice authentication on smart glasses remains challenging, as few studies have comprehensively explored acoustic features specifically suited to this form factor. For other wearable platforms such as smartphones and earables (zhang2016voicelive; li2020vocalprint; feng2017continuous; blue20182ma; meng2022your; gao2021voice), researchers have investigated a wide range of acoustic representations, including time-domain features (zhang2016voicelive), frequency-domain features (huang2025eve), time–frequency features (li2020vocalprint; li2024boneauth), and sound-field features (meng2022your; blue20182ma). While these approaches have demonstrated high accuracy, their direct transfer to smart glasses is non-trivial due to the unique form factor and acoustic constraints of eyewear. In particular, smart glasses face challenges such as limited flexibility in microphone placement, variable distance from the user’s mouth, and increased susceptibility to ambient noise. All of these challenges hinder the straightforward adaptation of existing acoustic authentication techniques. Motivated by these challenges, we investigate features derived from both air- and bone-conducted acoustic domains, and evaluate their effectiveness in supporting robust voice authentication for smart glasses.
@@ 2.3. Liveness Detection for Voice Authentication
Beyond distinguishing users from impostors, voice authentication must verify that commands are genuinely produced by the user, a process referred to as liveness detection(robustDetect; li2024boneauth; akhtar2014liveness; voicelivenessdetection; zhang2017hearing; zhang2022livoauth). To achieve this, researchers have proposed various methods, which can be categorized into speaker detection and speech detection. Speaker detection methods aim to determine whether a given sound originates from an artificial source, typically by exploiting features associated with machine-based speech production mechanisms, such as acoustic features(ahmed2020void; robustDetect) and magnetic features(chen2017you). However, these methods are tailored to the specific design characteristics of certain artificial speakers, which inherently limits their generalizability. As speech synthesis and voice conversion technologies continue to advance, such approaches may face increasing risks of being circumvented, thereby undermining their robustness in practical deployment.
Speech detection methods, in contrast, exploit features directly associated with the speaking process to ensure that the captured signal is temporally aligned with the user’s actual speech production (ContinuousArti; sahidullah2017robust; li2024boneauth; zhang2024safari). Building on this idea, researchers have explored a range of approaches, including throat vibrations (sahidullah2017robust), bone-conducted vibrations (li2024boneauth), facial muscle movements (zhang2024safari), airflow during articulation (wang2019secure), and pop sounds (voicepop), among others. While these studies have achieved promising results, most of them depend on sensors that either require close contact with the skin (huang2025eve) or must be placed at locations impractical for everyday wearable devices (e.g., the throat) (sahidullah2017robust). Such constraints not only affect user comfort and social acceptability but also hinder seamless integration into mainstream wearables, particularly smart glasses. To address these challenges and enable practical voice authentication on smart glasses, we further investigate the feasibility of leveraging combined acoustic features for liveness detection.
@@ 3. Understanding Air-Bone Acoustic Features by Smart Glasses
Figure 1. (a) Structure of vocal tract, including vocal cords, tongue and other articulatory organs (b) Previous research revealed that phoneme sounds could be localized by method of Time Difference by Arrival (zhang2016voicelive). (c) The possible acoustic features related to speech production, including vibration conducted by skin and bones, time delay and energy decrease of sound field in spatial dimension.
@@ 3.1. Principle and Possibility
To understand how smart glasses can leverage voice-related features for authentication, it is essential to briefly review the human speech production system, as illustrated in Fig. 3a. Airflow from the lungs excites the vocal folds, the vocal tract then shapes this excitation into an acoustic realization with characteristic amplitudes and spectra (badin2006three). After that, through precise control of the tongue and muscles, humans can articulate different phonemes and ultimately form words and sentences. Crucially, speaker-specific anatomy yields distinct propagation paths, resulting in both post-utterance signals and corresponding transmission paths carrying user-unique signatures for authentication.
However, existing approaches are based on features extracted from either the post-utterance voice or the transmission path, limiting comprehensive usage of the voice generation process. Most researchers focus on post-utterance features including airflow (wang2019secure), pop-sound (voicepop), and knowledge/content-based cues built from the observed speech signal. In contrast, motion-based methods that leverage IMU sensors in smart glasses exploit propagation-induced vibrations of the head/face but have limited access to the linguistic/ phonic content itself—and in smart glasses(zhang2024safari). We therefore ask: how to integrate voice features with transmission-path cues on smart glasses to enable more robust authentication?
In this paper, we found smart glasses offer two potential sensing avenues that jointly encode voice features and how it propagates (Fig. 3b–c). First, air-conduction (air) microphone arrays in the frame enable capturing both the voice and the path-dependent spatial features of the surrounding sound field shaped by head geometry and motion of the articulator (zhang2016voicelive). A bone/skin-conduction microphone (bone) at a stable contact point (e.g. nose pads) senses the voiceprint plus the propagation-induced vibrations transmitted through tissue and bone (Vocal_resonance). Thus, both channels inherently provide comprehensive features along the voice generation process, potentially enabling more robust authentication for smartglasses.
@@ 3.2.1. Design and Implementation
To capture the sound field as accurately and comprehensively as possible, we arranged a total of 14 omnidirectional air-conduction (AC) microphones around the user’s head (infineon IM73D122V01(Infineon)) and oriented most of them downward. To maximize generality, our placement strategy spans multiple commercial form factors — including Ray-Ban Meta (nose bridge and legs), Bose Frames (behind the ears). We also integrated microphones on the rims as some prior work also investigated acoustic-based interaction and health sensing at the same locations (echobreath; Echotouch).
In order to capture vibration features, we employed bone-conduction microphones (VPU14AA01 BC)(Sonion) for data collection, as they provide reliable sensing when in direct contact with the skin. Since smart glasses only make stable contact with the head around the nose and ears, and pilot studies showed that microphones near the ears were easily affected by hair friction, we mounted the bone-conduction microphones on each side of the nose pads to ensure better recording quality. The overall microphone layout is illustrated in Fig. 2a.
To acquire high-fidelity acoustic signals and extract meaningful insights, both bone-conduction and air-conduction microphones were configured in synchronization with a 96 kHz sampling rate. To achieve this, a collection board was used to connect all microphones simultaneously via flexible flat cables (FFCs), as shown in Fig.2a. The collection board was driven by an Orange Pi(OrangePi), which also transmitted data to a PC over a network cable using the TCP protocol. Note that a wire bundle was designed to prevent crosstalk between adjacent FFCs and to keep the cables organized, while maintaining user comfort.
@@ 3.2.2. Data Collection
Figure 3. Raw data from AC BC microphone and their Mel spectrum with phoneme segmentation by MAUS.
Figure 4. Mel-spectrograms of acoustic signals collected from different users using AC and BC microphones.
3.3. Key Findings and Insights
After data collection, we then analyzed the acoustic features of speech. Based on the mechanisms of human speech production, we examined features that capture the unique characteristics of individuals when speaking by vibration characteristics and sound field features.
3.3.1. Acoustic Vibration Features
Using the bone-conduction (BC) microphones embedded in the nose pads, we collected two channels of vibration signals. As some users maintained close contact with only one side of the nose pad, we selectively utilized the channel that yielded the more stable signal. Previous studies have demonstrated a correlation between signals captured by bone-conduction microphones and air-conduction microphones on earphones(li2024boneauth). Motivated by this observation, we compared the BC microphone output with that of the channel 7 (Fig 2 a) AC microphone, as it was located most closely to user’s mouth and captured most detailed signal. Fig.3 demonstrate a comparison of Mel-spectrogram of AC and BC signals. In line with prior work, the Mel-spectrograms of these two kind of microphones exhibited relative consistency during speech, and both of them successfully aligned with phoneme boundaries processed by Munich Automatic Segmentation System (MAUS(Maus)). We also compared the outputs of BC and AC microphones when different users spoke the same sentence, as shown in Fig.4. While both BC and AC signals were able to capture speech, they exhibited noticeable differences above 4kHzkHz. Therefore, we regarded them as distinct and usable features.
3.3.2. Sound Field Features
Figure 5. (a) Waveform differences between the reference AC microphone (channel 7) and other AC microphones, with amplitude differences and time delays annotated across channels. (b) The average time delay and energy ratio (using channel 7 as reference). The shaded area indicates the variation range. (c) Energy delay matrix (EDM) and time delay matrix (TDM) for a speech segment.
Figure 6. TDM asymmetry and Mel-spectrograms of speech segments from different users. In the TDM asymmetry figure, blue indicates that the time delay on the right side while wearing the smart glasses is greater than that on the left, whereas yellow indicates the opposite.
As previous study revealed that the time delay (zhang2016voicelive) and energy difference (yan2019catcher) of a certain sound wave could act as distinctive features for authentication, we tried to explore the sound field feature through these two aspects. Fig. 5a demonstrates comparison of audio data on different AC channels during a certain fraction of time. Since the AC microphone located above the nose and oriented downward (channel 7 in Fig 2) received the strongest energy, we utilized it as the reference channel and compare it with channels alongside one of the rim respectively. We observed that for channels along the rim, the relative audio energy with channel 7 decreased as the distance between the microphone and the mouth increased, while the time delay increased at the same time. Considering that the sound field may vary during people speaking, we detailed the changes by segmenting the audio into pieces, and calculated the energy ratio and time delay of all channels during each frame, forming an Energy Decrease Matrix (EDM) and a Time Delay Matrix (TDM), as shown in Fig 5c. From the figure, we noticed that the energy ratio and time delay captured by microphones at the position fluctuates continuously, which speaks to the facial movement of users.
Also, by measuring the changes of EDM and TDM alongside the time axis, we found that the maximum and minimum value of symmetrical pairs of AC microphones have noticeable difference, which speaks to the subtle asymmetry when capturing audios. To further evaluated this asymmetry, we calculated the time delay and energy ratio difference of symmetrical pairs by different users. Fig 6 demonstrated that different users produce subtle asymmetry patterns even when speaking the same content, making TDM and EDM potential features for authentication and anti-spoofing.